### v3

# Vessel Detection and Localization Using Distributed Acoustic Sensing in Submarine Optical Fiber Cables

Abstract
Submarine cables play a critical role in global internet connectivity, energy transmission, and communication but remain vulnerable to accidental damage and sabotage. Recent incidents in the Baltic Sea highlighted the need for enhanced monitoring to protect this vital infrastructure. Traditional vessel detection methods, such as synthetic aperture radar, video surveillance, and multispectral satellite imagery, face limitations in real-time processing, adverse weather conditions, and coverage range. This paper explores Distributed Acoustic Sensing (DAS) as an alternative by repurposing submarine telecommunication cables as large-scale acoustic sensor arrays. DAS offers continuous real-time monitoring, operates independently of cooperative systems like the Automatic Identification System (AIS), being largely unaffected by lighting or weather conditions. However, existing research on DAS for vessel tracking is limited in scale and lacks validation under real-world conditions. To address these gaps, a general and systematic methodology is presented for vessel detection and distance estimation using DAS. Advanced machine learning models are applied to improve detection and localization accuracy in dynamic maritime environments. The approach is evaluated over a continuous ten-day period, covering diverse ship and operational conditions, representing one of the largest-scale DAS-based vessel monitoring studies to date, and for which we release the full evaluation dataset. Results demonstrate DAS as a practical tool for maritime surveillance, with an overall $F_{1}{}$-score of over $90\%$ in vessel detection, and a mean average error of 141m141\,m for vessel distance estimation, bridging the gap between experimental research and real-world deployment.
I Introduction
Submarine cables, essential for global internet connectivity, energy transmission, and communication, are particularly vulnerable to accidental damage (e.g., fishing nets, ship anchors) and sabotage linked to geopolitical tensions [euromaidan_baltic_2025]. Recent incidents highlight this risk: in November 2024, damage to two Baltic Sea cables disrupted internet services between Finland, Germany, and Lithuania, with suspicions of foreign interference [cbsnews_baltic_cables_2024]. In December 2024 the EstLink2 power cable between Finland and Estonia was severed, prompting an investigation [navalnews_estlink2_2024]. In February 2025, the C-Lion1 cable between Finland and Germany suffered damage near Gotland Island [euromaidan_baltic_2025]. The European Commission recently addressed the need for stronger monitoring and security measures in its Joint Communication on submarine cable resilience [eu_submarine_cables2025].
Detecting and localizing vessels near submarine cables is thus crucial to prevent disruptions in international data communications. Existing vessel localization methods use various sensing techniques, each with strengths and limitations. Yu et al. [Yu_2021] demonstrated vessel detection with synthetic aperture radar (SAR), which enables wide-area surveillance in all weather conditions. However, SAR struggles with real-time processing and false alarms from sea clutter. Wawrzyniak et al. [Wawrzyniak_2019] used video surveillance in ports, applying vessel detection algorithms to track ships, but these systems are limited by lighting conditions and range. Similarly, Xie et al. [Xie_2020] improved ship detection in multispectral satellite imagery with machine learning (ML), enhancing performance under mist and clouds, though severe weather and nighttime conditions significantly reduce accuracy.
Distributed Acoustic Sensing (DAS) has emerged as a transformative technology enabling dense spatial sampling of acoustic signals along fiber-optic cables, and has seen increasing applications across geophysics, infrastructure monitoring, and environmental sensing. Recent reviews such as [cheng2024photonic] provide an overview of the evolution, challenges, and potentials of DAS across different domains. In the paper context, DAS overcomes many limitations of traditional methods by repurposing submarine telecommunication cables as virtual sensor arrays [Landro_2022]. Unlike SAR and optical imagery, DAS enables continuous real-time monitoring with high spatial and temporal resolution, independently of lighting and weather conditions [Thiem_2023]. It also operates without cooperative systems like the Automatic Identification System (AIS), allowing the detection of “dark” ships involved in unauthorized activities [Malaprade2019]. By making use of existing infrastructure, DAS provides wide-area coverage at lower costs without requiring dedicated hardware like hydrophones [Malaprade2019, Landro_2022]. DAS detects ships by analyzing their acoustic signatures, such as broadband noise from machinery and tonal signals from propellers. These characteristics have been studied for vessel classification, speed estimation, and maritime traffic monitoring [Malinowski_2002]. However, existing methods must be adapted to the specific signal-processing requirements of DAS. Despite advancements, significant gaps remain in the DAS literature for vessel detection and localization, specially those oriented to submarine cable protection. Many studies use limited datasets, lack diversity in vessel types and environmental conditions, or focus on proof-of-concept setups without real-world validation.
The main contributions of this work are:
  • A general and systematic methodology for vessel detection and distance estimation, applying DAS over submarine cables.
  • Use of advanced ML models including XGBoost and neural networks, to enhance vessel detection and localization in dynamic maritime environments.
  • A large-scale rigorous evaluation over real-world conditions, covering a continuous ten-day period with diverse ship types, sizes, and operational scenarios, far exceeding previous studies addressing vessel detection and localization tasks, for submarine cable protection applications.
  • The fully annotated dataset we used, which is available in a public repository.
  • Evidence of DAS as a practical tool for maritime surveillance, bridging the gap between experimental research and real-world deployment.
To the best of our knowledge, this is the first study to systematically address vessel detection and distance estimation using machine learning in submarine cable DAS data, specifically focused in a submarine cable protection task, at a much larger scale than any other previous work.
The remainder of this paper is structured as follows. Section II introduces the previous work on vessel detection, classification and localization using DAS. Section III provides a general overview of DAS technology in submarine cables, with Section IV describing the data used in this study. Section V provides details on the proposed system architecture and methodological aspects, while Section VI presents the experimental work, and the obtained results. Finally, Section VII concludes the paper and outlines future research directions.
II Previous Work
DAS data present significant challenges due to its high-dimensionality, noise levels, and large volume, often surpassing the capabilities of traditional data storage and signal processing capabilities [Nur_DAS_storage2024, dong2020dassa, Rivet_2021]. On the other hand, advanced techniques for data interpretation such as artificial intelligence and ML methods have demonstrated potential in overcoming these challenges for vessel detection [Malaprade2019, Rivet_2021], seismic monitoring [Stork_2020], and acoustic recognition [Drylerakis2024source, zhan2024application]. Malaprade et al.[Malaprade2019] illustrated the feasibility of ML-based ship detection using frequency band energy plots but were limited by a small dataset (only seven crossings) and sparse experimental details. Thiem et al.[Thiem_2023] localized a single passenger ship in a Norwegian fjord using noise-reduction filtering and traveltime inversion, showing good agreement with AIS but focusing on just one vessel type and not assessing other factors like tides or weather. Rivet et al.[Rivet_2021] estimated tanker trajectories off Toulon, France, leveraging Doppler effects and simulation models; while results aligned with observations, the study considered only one vessel and lacked broader environmental variability. Chen et al.[Chen_2022] deployed a photonic integrated sensing and communication system in the Pearl River estuary, effectively tracking a ferry route, though only one vessel type and route were examined. Wienecke et al.[wienecke2023new] detected bottom trawlers via seabed vibrations up to 2.5km2.5\,km from the cable, but tested no multi-vessel scenarios. Similarly, Drylerakis et al.[Drylerakis2024source] used PCA-based denoising to improve signals from a single passing vessel.
Other works emphasize shorter cable deployments or more constrained tests. Dias et al.[dias2014] used a 500m500\,m cable in shallow waters to detect artificial ship noise, while Douglass et al.[Douglass_2023] captured broadband acoustic signals along a 3.5km3.5\,km long cable and compared DAS responses to hydrophones. Paap et al.[paap2025leveragin] demonstrated automatic detection and localization for three specific vessels, achieving good performance but not exploring more diverse vessel types. Shao et al.[shao2025tracking] used Doppler shifts to track ships in a short 180m180\,m river cable deployment; the method aligned well with GPS but only covered three vessel passages.
Very recently, Huang et al.[Huang_2025] proposed DASHip, a large-scale labeled dataset for ship detection based on AIS-guided annotation of 55,875 events. They reported strong detection performance (73.6% joint detection ratio) using an object visual detection algorithm (YOLO), yet important higher-frequency acoustic cues (10200Hz10--200\,Hz) were sacrificed by downsampling to 10Hz10\,Hz, and their wake-centric approach could not identify vessels before cable crossings. This domain gap between clean labels and real-world conditions potentially reduces generalization[Karasalo_2017, Landro_2022].
Beyond direct ship detection, Landrø et al.[Landro_2022] analyzed various acoustic phenomena, including seismic events and whale calls, using an Arctic cable. They tracked an 86-meter cargo ship via hyperbolic wavefront analysis and near-field beamforming. Stork et al.[Stork_2020] likewise demonstrated microseismic and whale vocalization detection on a 2km2\,km cable, highlighting the versatility of DAS for multi-purpose acoustic monitoring. While such studies do not center on vessel detection, they confirm DAS’s broad applicability in ocean acoustic sensing.
From the above discussion, despite demonstrating the feasibility of DAS-based vessel detection, previous works tend to rely on very small or narrowly focused datasets, single vessels, or limited scenarios. Many studies lack broad environmental variability, often considering short cable segments, and do not systematically address cable protection goals such as quantifying vessel proximity or implementing detection prior to a direct cable crossing. In contrast, our proposal emphasizes explicit cable protection objectives by detecting vessel presence and estimating distance under diverse environmental conditions, on a substantially larger dataset than typically reported in the literature.
(a) Regional map.
(b) Local map showing cable (red line).
(c) Local map showing cable (red line) and bathymetry.
Figure 1: General location and bathymetry (cable location has been displaced for security considerations).
III Distributed Acoustic Sensing in Submarine Cables
Optical fiber-based distributed acoustic sensors are advanced sensing systems that use optical fibers as the primary medium for detecting and monitoring dynamic acoustic signals along their entire length. Unlike traditional sensor networks, DAS systems convert a conventional optical fiber into a vast array of virtual sensors that can measure dynamic strain changes along its entire length [Zhan_2020]. Conventional DAS technology probes the fiber-optic cable with a coherent laser pulse, and measures changes in the phase of the returning Rayleigh back-scatter signal time-series. Optical phase shifts in the back-scattered signal are proportional to longitudinal deformation in the fiber and can be translated into local strain changes across a fiber segment (termed gauge length $L$) by integration. The primary issue with the conventional phase-measuring approach is that a substantial fraction of positions along the fiber exhibit zero or near-zero intensity. These low-intensity points (typically called “fading points”) in the back-scattered signal lead to a defective determination of the phase of the local back-scattered field, implying inhomogeneous sensitivity variations across the DAS array [gabai2016sensitivity]. Phase-measuring DAS systems generally available in the market achieve a fully distributed measurement of vibrations of up to several kHz, with a resolution of a few meters and measurement ranges of a few tens of kilometers. Optical coding methods are one of the preferred approaches for extending the measurement range of DAS systems. Among these, optical pulse compression reflectometry [Zou_2015] has been widely developed, allowing measurement ranges well exceeding 100km100\,km [Waagaard_2021].
Regardless of the interrogation method, DAS sensitivity depends on the optical fiber type, installation conditions, and local fiber-ground coupling. In existing installations, these factors are essentially not controllable; thus, effective data processing algorithms must account for them. Recent advances in signal processing and machine learning techniques have enabled improved handling of coupling variations, significantly enhancing DAS performance and making it a powerful tool for complex, large-scale monitoring tasks [munoz2022enhancing].
One of the mainstream applications of DAS is the monitoring of submarine cables for various purposes. In particular, DAS technology has proven effective in the surveillance of both communication and energy transmission submarine cables [Brenne2019, Malaprade2019]. Traditional offshore monitoring methods typically rely on fixed platforms, buoys, or ocean-bottom point sensors, offering limited spatial coverage. This constraint hinders the detection of dynamic and spatially diverse events, such as maritime activity involving vessels, large marine mammals, or other wildlife, occurring at the sea surface or seabed.
IV Database Description
IV-A Geographical and Instrumentation Details
In our study, we used DAS and AIS data collected over a 10-day period between June $16^{th}$ and $25^{th}$ 2023, to develop models for vessel detection and localization.
The dataset was recorded using a 28km28\,km long pre-existing ocean-bottom fiber-optic cable in the Southern Bight of the North Sea offshore Zeebrugge, Belgium (see Fig. 1). The fiber-optic cable was originally installed to monitor a power cable from an offshore wind farm facility. The cable is buried between 1.4m1.4\,m and 7.2m7.2\,m below the seafloor, with an average burial depth of 4.7m4.7\,m, which provides a good acoustic coupling between the seabed vibrations and the fiber [paap2025leveragin]. The Alcatel OptoDAS interrogator [OptoDAS] was used for strain data acquisition. It is a phase-measuring DAS system that uses optical pulse compression reflectometry to extend the measurement range. The gauge length in this case is $L\approx 10\,m$, which is significantly larger than the nominal optical resolution of the system (1m1\,m), related to the bandwidth of the optical probing signal used. This significantly reduces the sensitivity issues related to fading points.
The interrogator operates with a spatial resolution of 10.21m10.21\,m (channel spacing), creating $2774$ simultaneously recording strain sensors, and generating differential phase data signals at a $f_{s}=3125\,Hz$ sampling frequency. These raw signals are further preprocessed following the procedure described in Section V-A.
IV-B Metadata Processing
Effective analysis of the DAS data required careful processing and integration of several crucial auxiliary datasets (metadata). This section details the acquisition, characteristics, and processing steps applied to the vessel tracking data (AIS), the fiber’s geographical and bathymetric information, and addresses key considerations regarding data balance and spatio-temporal calibration necessary for the subsequent analysis. Fig. 6 illustrates how these metadata processing steps are integrated within the overall system architecture presented in section V.
IV-B1 AIS Data
In this study, AIS data were employed primarily for geographically labeling the DAS data, providing ground-truth vessel positions, which are essential for both training and evaluation purposes [emmens2021promisesAIS]. The raw AIS data, provided by the cable owner, included Maritime Mobile Service Identity (MMSI) numbers, UTC timestamps, vessel positions (longitude and latitude), course, speed, and heading for 745 unique vessels, totaling 64,417 reported positions. We further augmented this dataset by web-scraping static vessel information (name, IMO number, type, dimensions, tonnage), although dynamic details like navigational status and draught were unavailable. Analysis of vessel types revealed 45 distinct categories, with cargo ships being the most prevalent in the monitored area, alongside notable presences of fishing vessels and tugs (see Table I for detailed statistics).
TABLE I: Vessel type distribution in AIS data (sorted by occurrences).
# vessels % total Avg length
Container Ship 134 18% 263
Chemical/Oil Products Tanker 124 17% 151
General Cargo Ship 96 13% 114
Vehicles Carrier 62 8% 182
Bulk Carrier 61 8% 194
Unknown 41 6% N/A
Ro-Ro Cargo Ship 39 5% 201
Lpg Tanker 29 4% 128
Hopper Dredger 19 3% 106
Fishing Vessel 19 3% 22
Tug 16 2% 31
Crude Oil Tanker 10 1% 248
A significant challenge encountered when utilizing AIS data for machine learning applications, including ours, is the typically low and variable update rate [Harati-Mokhtari_Wall_Brooks_Wang_2007AIS, emmens2021promisesAIS]. While ITU recommendations suggest reporting intervals under 30 seconds for most moving vessels [ITU_R_M1371_5AIS], our dataset exhibited much lower frequencies, with $88\%$ of vessels reporting only every 1 to 3 minutes (see Fig. 2 for details). Considering the average vessel speed of 10.2 knots observed in the dataset (Fig. 3), such infrequent updates can lead to significant positional uncertainty (approximately 315 meters per minute in average), preventing accurate labeling. To address this limitation, we applied linear interpolation between reported positions with a 1-second resolution, excluding intervals with vessels not reporting during 60 minutes, and incorporating additional heuristic rules to minimize the generation of erroneous trajectories. Given the update rate distribution and the good results achieved, we believe the linear interpolation is good enough for the explored scenario, although in more demanding situations, alternative interpolation methods should be adopted [guo2021improved] (more details in the supplementary material Web page). This interpolation process resulted in a significantly denser dataset of 1,207,446 position entries used for subsequent analysis.
Figure 2: AIS data update rate histogram for the used dataset.
Figure 3: AIS vessel speed histogram for the used dataset.
IV-B2 Geometrical + bathymetry data acquisition
Accurate knowledge of the fiber-optic cable’s geographical position and the surrounding bathymetry is critical for DAS-based location-dependent monitoring. Variations in seafloor morphology and cable burial conditions directly impact the coupling between the cable and the seabed, significantly influencing the quality and characteristics of the recorded acoustic signals [sladen2019distributed, Mata_Flores_2023]. Furthermore, precise fiber localization is essential for accurately positioning detected events, such as vessel passages. While cable deployment aims to follow a specific trajectory, the final resting position on the seafloor can deviate, and its burial status may change over time [sladen2019distributed].
For this study, we utilized precise geometrical and bathymetry measurements for the full 28km28\,km cable length, provided directly by the cable owner. This dataset offers significantly higher precision than publicly available sources like EMODNet [emodnet2024] (which has a resolution of 115 meters). A comparison, shown in Fig. 4, reveals depth discrepancies of up to 6 meters between the owner’s data and EMODNet along the cable path used in this work.
Fig. 4.a illustrates the bathymetry along the entire cable, while Fig. 4.b provides a detailed view of the specific segment used for our experiments. This region corresponds to a section where the cable crosses a dredged artificial channel (approximately 20m20\,m deep) providing harbor access. Here, the cable itself is buried at depths reaching 25m25\,m. Our experiments analyzed data from segments spanning 102.1m102.1\,m to 2552.5m2552.5\,m (corresponding to $10$ and $250$ sensed positions), as indicated by the highlighted blue and pink regions in Fig. 4.a and Fig. 4.b, respectively. The use of this high-resolution, owner-provided data ensures accurate geographical referencing for our analysis.
(a) Full fiber-optic cable.
(b) Selected fiber-optic cable region.
Figure 4: Comparison of bathymetry sources in the cable under study. Sea floor depth provided by the cable owner (orange trace) and EMODnet (green dotted line). The Fiber position (buried cable) is also shown (gray trace). Diff. EMODnet vs. owner measures depth differences between EMODnet and owner data (dark blue trace). Selected fiber-optic cable regions are shown in blue background (250 sensors range) and pink background (50 sensors range).
IV-B3 Data calibration and synchronization
Accurate geographical localization is essential, requiring each DAS sensed data point to be mapped first to a specific location along the fiber, and then to its corresponding geographical coordinates (longitude, latitude, and depth). This geographical calibration process typically relies on identifying anchor points (locations with known coordinates) along the fiber. Establishing the mapping between fiber distance and geographical position, whether automatically or semi-automatically, can be challenging, particularly for sections far from the sensing equipment.
The best approach to fulfill the geographical calibration task would be direct measurements of the cable position (longitude, latitude and depth) combined with knowledge on anchor points of well-known fiber positions. This scenario is not typically realistic, so that alternative methods based on general bathymetry sources combined with cable deployment information and interpolation strategies can also be used.
Finally, the availability of UTC timestamps in AIS data allows time synchronization with the acquired DAS data, which is also accurately timed.
IV-C Data Balance Considerations
The classification task categorizes data frames based on vessel proximity relative to a distance threshold. Our dataset exhibits significant class imbalance, meaning the number of frames with vessels closer versus further than the threshold is often unequal. This imbalance varies considerably both along the fiber’s length and depending on the specific distance threshold chosen. As an example, Fig. 5 shows the distribution along the fiber length (horizontal axis) of the number of available data frames with a vessel closer than 1000m1000\,m (blue area), data frames with vessels further than 3000m3000\,m (reddish area), and data frames with vessels at distances between them (orange region). Consequently, these data balance characteristics were crucial considerations in our experimental design and required careful interpretation of results, particularly favouring performance scores that properly assess the results in unbalanced scenarios.
Figure 5: Class instance distribution for different distance thresholds.
IV-D Data Availability
Due to data-owner restrictions, neither the original raw differential-strain recordings nor precise geographical localization can be shared publicly. Instead, we have released the Marlinks-NS DAS dataset in Zenodo [ramirez2024dasvesseldataset], that includes the complete set of processed data for the experiments in this work. It comprises the feature vectors and their corresponding ground-truth labels (timestamp, closest-vessel distance information, and vessel related metadata) in a structured hdf5 archive. To ease data distribution and reproducibility, we have also generated a GitHub repository (available at https://github.com/UAH-PSI/das-vessel-detection) that provides the data description, initial source code for loading and processing the data, as well as a small showcase data subset.
V DAS+ML system for vessel detection and localization
The overall system architecture proposed in this work is shown in Fig. 6, which we detail next.
Figure 6: Proposed System Architecture (with reference to the corresponding paper section in relevant modules).
V-A DAS Signal Preprocessing
The DAS interrogator used in this work is a phase DAS, and the generated raw signal is differential phase. To convert this into a direct strain measurement, the following signal preprocessing operations are carried out:
  • Data scaling: Adjusts the raw differential phase measurements based on system gain and calibration coefficients, ensuring subsequent steps operate on signals with consistent amplitude levels.
  • Spike detection and removal: Removes impulsive spikes using an adaptive thresholding method, where samples exceeding a sliding-window mean plus an offset are set to zero before integration.
  • Phase unwrapping: Converts the wrapped phase (limited to $\pi,\pi)$) into a continuous time series by correcting abrupt $2\pi$ jumps, thus preserving the true optical path length changes.
  • Time integration: Integrates the unwrapped phase over time and scale it to recover the cumulative strain in the fiber. This step yields a direct strain signal rather than a differential measurement. Fig. 8.a shows an example of a strain signal acquired during a vessel crossing above the cable.
Strain-rate and strain are linearly related in the frequency domain (essentially involving a multiplication of the strain-rate spectrum by $1/j\omega)$, so that the use of either magnitude would simply lead to changes in model tuning, and being related by a linear operation, it should definitely not affect the system performance.
V-B Spectral Feature Extraction
Extracting robust features from DAS data for vessel detection and localization is complicated due to the intricate physical processes converting ship noise into measured strain. Accurately modeling this conversion is challenging due to several factors: the complex and variable nature of noise radiated by different vessels [Rivet_2021]; the complexities of underwater acoustic propagation, including attenuation, reflection, and Doppler effects [Karasalo_2017, vadov2001long, li2024precise]; variations in the coupling between the fiber cable and seabed, which strongly influence strain measurements [Lior_2021, Mata_Flores_2023]; and potential artifacts introduced by the DAS system itself (due to electronic and optical design, and the conditions in the interrogator deployment environment).
Given that acquiring the precise knowledge needed for a full physics-based model is often impractical due to environmental variability and incomplete deployment information, our proposal is adopting a purely data-driven strategy. A frequent technique involves analyzing the energy within specific frequency bands, which has proven effective for interpreting DAS signals and providing features for associated ML tasks [tejedor2016toward, hartog2017introduction, Malaprade2019]. Adopting this principle, our work utilizes a data-driven spectral analysis, informed by available AIS data, to generate feature vectors suitable for guiding the ML process. Further details are given in Section VI-C, as it implies experimental work for feature vector design.
V-C Machine Learning Tasks and Training Modules
The experimental work can be oriented to different ML tasks. In this work, we focus on vessel detection and localization, as follows:
  • Vessel detection: This addresses a binary classification task, and aims to detect the presence of any vessel at a distance closer than a given threshold. This threshold must be designed considering criteria related to security, fiber response capabilities, and data balance. Regarding security considerations, in a real world deployment, an alarm should be triggered if the vessel is a potential threat, and with enough time in advance to be able to take the necessary measures. As an example, a Spanish company managing deployed submarine cables being monitored with AIS data to prevent vessel threats to the cable, sets a distance threshold of 1000m1000\,m to consider further detailed vessel tracking. Also, a suitable distance threshold should be chosen to allow a detectable response in the fiber (which may be an issue in case buried cables are used), with a reasonable data availability balance between the two classes (threat and non-threat).
  • Vessel localization: This addresses a regression task, aimed to estimate the distance of the closest vessel to the cable. In this case, distance thresholds need to be set to deal with distance estimation at different distance ranges, assuming that different models will be built for each range.
To tackle these detection and localization tasks, we evaluated two well-established machine learning approaches:
  • XGBoost (eXtreme Gradient Boosting): An optimized gradient boosting framework that learns an ensemble of decision trees [Chen:2016:XST:2939672.2939785]. XGBoost was selected for its proven effectiveness, scalability to large datasets, and inherent mechanisms to mitigate overfitting, making it suitable for handling potentially high-dimensional feature spaces derived from DAS data.
  • Neural Network (NN): A simple feed-forward neural network architecture was also implemented, providing a contrast based on NN learning principles.
In both classification and regression scenarios, an associated training process is required to generate the corresponding models. The availability of sufficient and adequately balanced data among the considered classes or distance ranges is crucial. This issue can be particularly significant in marine environments around deployed submarine cables, as maritime traffic varies heavily. Therefore, a detailed analysis of the available data is necessary to assess its suitability and guide the data acquisition and partitioning tasks.
Another key element that impacts the reliability and robustness of a system and its real generalization capabilities is the rigor of the experimental approach [tejedor2017machine]. In time-series data, such as DAS recordings, temporal dependencies can introduce biases if not properly handled. The typically used random partitioning strategy may result in training and validation sets containing temporally adjacent data, leading to information leakage and overestimation of model performance. Special care must be taken in this stage, adopting specific cross-validation strategies that include restrictions on the selection of temporal ranges for training and testing subsets.
Further details are given in Section VI-D.
V-D Result Post-Processing
In ML tasks dealing with time series, we can further improve the classification results by applying post-processing methods. Typical strategies are smoothing via sliding window averaging, median or mode filtering, majority voting (using ensembles or temporal ranges), multiple intervals, etc. [bagnall2017great].
Finally, one key benefit of DAS is the extensive range of sensed positions, which allows the systems to exploit the spatial and temporal redundancy of the acquired data. This should be explicitly included in either the feature extraction or ML stages. Suitable alternatives include exploiting array signal processing [munoz2022enhancing], or data integration across temporal and spatial ranges [tejedor2019contextual, huynh2025real], which will be detailed in Section VI-D.
VI Experimental Work, Results and Discussion
VI-A Database Partitioning
Given the characteristics of our task, we adopted the $k$-fold cross-validation strategy, a widely used technique for assessing model generalization capability [Wong2019-kfold]. This involves dividing the whole dataset into $k$ equal subsets (folds). The model is trained on k1k-1 folds and validated (tested) on the remaining fold. This process repeats $k$ times, with each fold serving once as the test set. The results from all test sets are then averaged to estimate the overall performance. Therefore, this method ensures that every data point is used for both training and testing, providing a comprehensive evaluation of the model’s performance while maintaining a strict separation between training and testing data.
As described in Section IV, the dataset was collected over a 10-day period to develop models for vessel detection and distance estimation. Therefore, to ensure a rigorous and robust evaluation and mitigate potential biases, we adopted a 10-fold cross-validation strategy, assigning each day’s data to a distinct fold. That way, we maintain the temporal continuity and integrity within each fold and also ensure that the testing set comprises data points that are not temporally interleaved with those of the training set. Additionally, this data partition strategy ensures that higher variability is introduced in the model generation, promoting a better generalization on unseen data.
Considering all $k$-folds, it is worth mentioning than only 11 vessels (out of $565$ in the vicinity of the explored sensor range) appear in more than 5 folds along the recording period, which indicates the difficulty of the task.
VI-B Performance Metrics
The performance metrics depend on the addressed task (vessel detection or vessel localization). For vessel detection, we define Class 0Class\,0 as the condition in which a vessel is closer than a predefined threshold distance, and Class 1Class\,1 when there is no vessel closer than the predefined threshold distance. The following classification performance metrics are used:
  • Accuracy: This is defined as the percentage of correctly classified instances for all classes (the higher, the better).
[…]
    We report both the global $F_{1}$-score (referred to from now on as $F_{1}^{G}{}$) and class-wise $F_{1}{}$-scores (referred to as $F_{1}^{C0}{}$ and $F_{1}^{C1}{}$ for class 0 and class 1 $F_{1}{}$ scores, respectively).
On the other hand, for vessel localization, the regression model is evaluated using the Mean Absolute Error (MAE). This measures the average absolute deviation between predictions and actual values (the lower, the better):
$MAE=\frac{1}{N_{F}{}}\sum_{i=1}^{N_{F}{}}|y_{i}-\hat{y}_{i}|,$ (1)
where $N_{F}$ is the number of data frames for which we generate a distance estimation value $\hat{y}_{i}$, and $y_{i}$ is the real distance.
In the vessel distance-estimation task, where the MAE will be calculated on data subsets with different distance thresholds, we will also normalize the MAE by these distance thresholds ($\delta_{thr}\in D_{thr}$), to get a relative estimation of the error, as compared with this distance threshold: $rMAE=\penalty 10000\frac{MAE}{\delta_{thr}}$.
To compare the results between the two ML approaches, we will quantify the uncertainty of our performance metrics using a non-parametric bootstrap procedure [efron1994introduction] (adapted to our $k$-fold cross-validation strategy) that has been further validated by subsequent works on bootstrap methods in ML [goo2024confidence, shabbir2024estimation].
VI-C Feature Extraction Analysis
The initial spectral analysis we carried out to guide our feature extraction design (as discussed in Section V-B) is based on the calculation of long term spectral averages for different environmental conditions related to the presence or not of nearby vessels. We defined nearby in relation to different distance thresholds with respect to fiber sensing positions. Specifically, calculations were computed using data frames corresponding to fiber positions:
  • That had no nearby vessels, with distance thresholds of 3000m3000\,m and 5000m5000\,m (condition referred to as noise).
  • That had nearby vessels, with distance thresholds of 1000m1000\,m and 2000m2000\,m (condition referred to as vessel).
The computation of the spectrum average for each recording is carried out from $10$-second windows. We also limited the displayed bandwidth to 250Hz250\,Hz as we found no significant information to appear in higher frequencies.
(a) Spectral content in the (0.1(0.1-250)Hz250)\,Hz range.
(b) Zoomed spectral content (vessels only).
(c) Spectral content for feature vector design.
Figure 7: Long term averaged strain power spectrum plots for strain measurements under different environmental conditions (expressed in dBs).
Figure 8: (a) Raw strain signal sensed at 15.4km15.4\,km in the time interval of the red dashed square in Fig. 8.d. (b) Filtered strain signal (4Hz4\,Hz to 98Hz98\,Hz, excluding the (4951)Hz(49-51)\,Hz) corresponding to the raw signal in Fig. 8.a and time-synchronized with it. (c) Filtered strain data as a function of time and distance along the cable corresponding to the red dashed square in Fig. 8.d, time synchronized with Figs. 8.a and 8.b. (d) Feature vector energy (in dBs) as a function of time and distance along the cable. (e) Distance from cable to nearest vessel (in meters), time-synchronized with Fig. 8.d.
Fig. 7.a shows the long term averaged strain power spectrum (expressed in dBs) for all environmental conditions in the (0.01250)Hz(0.01-250)\,Hz range. From the calculated averaged spectra, it can be seen that:
  • There is a constant strong peak located at 100Hz100\,Hz, with clear harmonics and subharmonics at 50Hz50\,Hz, 200Hz200\,Hz (and 300Hz300\,Hz, 400Hz400\,Hz and 500Hz500\,Hz not displayed), with smaller but clear peaks between 100Hz100\,Hz and 200Hz200\,Hz (and between 300Hz300\,Hz and 400Hz400\,Hz not displayed). The fact that these components appear for both noise and vessel spectra, suggest that they are probably caused by physical or mechanical systems (motors, rotating shaft, vibrating structures, etc.) in the interrogator environment (Fig. 7.b shows vessel spectra only, to allow comparison of the harmonic peaks).
  • There is a clear distinction between the average spectrum for the noise and vessel conditions in the (2100)Hz(2-100)\,Hz, opening the path to derive discriminative features for ML tasks (see zoomed spectra in Fig. 7.b).
  • There seems to be no differences between the considered noise conditions (for 3000m3000\,m and 5000m5000\,m distance thresholds).
  • The impact of the distance threshold in the vessel average spectrum is consistent: the closer the distance, the stronger the amplitudes. It is also important to note that the spectral shape is maintained for the considered thresholds, so that the generated ML models will face similar spectral conditions.
These conclusions guided our feature vector design, which will be composed of energy values in logarithmically spaced frequency bands (as shown in the logarithmic frequency axis in Fig. 7), to adequately represent lower frequency content. Each vector component is calculated by summing the FFT spectral coefficients corresponding to each defined frequency band. The considered bandwidth ranges from 4Hz4\,Hz to 98Hz98\,Hz, excluding the (4951)Hz(49-51)\,Hz interval to remove the 50Hz50\,Hz and 100Hz100\,Hz peaks shown in the spectral analysis (see Fig. 7.c), so that they will not mask the events captured by the DAS system. Fig. 8.b shows an example of the filtered strain signal corresponding to the raw strain signal in Fig. 8.a. An example of the evolution of the filtered strain as a function of time and distance along the cable is shown in Fig. 8.c.
[…]
As a prospective study of the feasibility of the proposed ML tasks, we generated a number of visualizations to relate the extracted features with the target elements, from which Figs. 8.d+e show an example. Fig. 8.d displays the temporal variation (horizontal axis, spanning for two hours) of the feature vector energy measured at each sensor (vertical axis, corresponding to $250$ sensors, spanning for 2553m2553\,m, from $14702$ to $17255$ meters in fiber distance from the interrogator). The feature vector energy is calculated as the sum of the energy components in all the selected frequency bands, and getting its value in db. Fig. 8.e is time synchronized with Fig. 8.d and represents, at each time instant, the distance from the sensor range to the closest vessel, between 0 and 2km2\,km (higher distances are saturated at the dark blue constant color to resemble lack of detected activity in the fiber). From Figs. 8.d+e, we can clearly see a very good match between the fiber measurements and the presence of vessels as they approach the fiber, cross over it, and progressively move away. We can also see different magnitudes of the vessel effect on the fiber (with longer or shorter spatial and temporal impacts), and with a clear relationship with the vessel presence in all the cases. Visible fiber response is detected along the full displayed fiber distance range, and up to roughly 10 minutes before and after the vessel crossing instant, which at an average speed of 10.2knots10.2\,knots implies a distance above 3km3\,km.
VI-D Experimental Setup
The feature vector used in our experiments follows the considerations outlined in Section V-B, and the results discussed in Section VI-C. We computed $N_{P}=100$ logarithmically distributed energy band values spanning the 4Hz4\,Hz to 98Hz98\,Hz bandwidth, excluding the (4951)Hz(49-51)\,Hz interval. The window length for feature vector computation is 10 seconds. Thus, the total number of feature vectors (corresponding to processed data frames) is $74771$, which will be distributed in training and testing subsets using the $k$-fold cross-validation approach described in Section VI-A.
We evaluated experimental conditions based on specific distance thresholds separating the vessel presence (Class 0) and absence (Class 1) for the classification task. The set of thresholds used, measured in meters, was $D_{thr}\penalty 10000=\penalty 10000\{500,1000,1500,2000,3000,5000\}$. For the vessel localization (regression) task, separate models were trained to estimate distances within ranges derived from these thresholds.
We also explored algorithmic strategies exploiting spatial and temporal redundancy (graphically summarized in Fig. 9):
  • Spatial redundancy: With configurations using $N_{S}=\penalty 10000\{10,25,50,100,250\}$ sensors, approximately spanning $100$, $250$, $500$, $1000$, and $2500$ meters, respectively.
  • Temporal redundancy: With temporal contexts of $S_{C}=\penalty 10000\{10,30,50\}$ seconds, using feature vectors from 1, 3, or 5 consecutive 10-second windows.
  • Feature averaging: Averaging features across the spatial or temporal dimensions, or both. Spatial averaging is always carried out over the selected sensor range.
  • Majority voting: Applied majority voting schemes with window lengths $V_{C}=\{1,3,5\}$ (where 1 means no voting). In our implementation, we consider v windows with the evaluated ss window length using a temporal shift of 10 seconds. For example, applying majority voting with $\texttt{v}=5$ windows of $\texttt{ss}=50\,s$, the considered temporal span is $90$ seconds. We did not evaluate longer temporal spans nor majority voting windows to avoid excessive waiting time in a realistic deployment.
Figure 9: Summary of frame processing strategies to exploit spatial and temporal redundancy in the feature space, and majority voting in the output results.
The specific configurations for the chosen machine learning models were as follows:
  • XGBoost: The baseline model definition was used with the following key parameters:
    • Objective function: Binary cross-entropy for classification and mean squared error for regression.
    • Booster type: Gradient boosted trees (gbtree).
    • Learning rate (eta): $0.05$.
    • Maximum tree depth (max_depth): $10$.
    • Number of boosting rounds (n_estimators): $500$.
  • Neural Network (NN): A simple architecture comprising:
    • Three fully connected (dense) layers.
    • ReLU (Rectified Linear Unit) activation functions.
    • Dropout rate of $50\%$.
    • Input layer size matching the feature vector dimension (dependent on averaging strategy).
    • Output layer size of 2 (for classification probabilities) or 1 (for regression distance).
    • Training batch size of 32.
Figure 10: Baseline $F_{1}{}$-scores as a function of distance threshold (T[500,1000,1500,2000,3000,5000]_V1_S10_ACH), by varying the number of sensors used ($N_{S}=\penalty 10000\{10,25,50,100,250\}$): (a) $F_{1}^{G}$, (b) $F_{1}^{C0}$, (c) $F_{1}^{C1}$.
Figure 11: Baseline $F_{1}{}$-scores as a function of distance threshold (T[500,1000,1500]_V1_S1_ACH) by varying the number of sensors used (10, 25, 50, 100 and 250): (a) $F_{1}^{G}$, (b) $F_{1}^{C0}$, (c) $F_{1}^{C1}$.
To mitigate overfitting and improve generalization, early stopping was applied during model training for both XGBoost and the NN. A validation set was systematically extracted from each training fold and used to monitor performance. Training was stopped if the relevant evaluation metric (e.g., validation loss or $F_{1}{}$-score) did not improve for 50 consecutive iterations.
To provide a clear view on the algorithmic variations in the included figures, we will refer to the different variants with the following notation: Ttttt_Vv_Sss_Aaa, where tttt refers to the distance threshold ($\texttt{tttt}\in D_{thr}$), v will be the number of temporal windows for majority voting ($\texttt{v}\in V_{C}$), ss will be the considered window length in seconds ($\texttt{ss}\in S_{C}$), and aa states if the feature vector averaging is done on temporal and/or spatial dimensions: $\texttt{aa}=\texttt{CH}$ for spatial averaging, and $\texttt{aa}=\texttt{TI}$ for combined spatial and temporal averaging. Note that spatial averaging is always carried out, with different number of sensors.
VI-E Vessel Detection Experiments
We will describe here the experiments using the XGBoost system, to select the best algorithmic variants. Then, we will compare it with the NN approach.
For the classification task, we first evaluated the baseline strategy with no majority voting, and $10$ seconds temporal context, only exploiting spatial redundancy using the considered number of sensors $N_{S}$.
Fig. 10 shows the $F_{1}{}$-scores for different number of sensors (horizontal axis), and different distance thresholds. The first observation is the relatively high performance rates, most of them being above $80\%$, with peaks above $90\%$. Considering the $F_{1}^{G}{}$-score (Fig. 10.a), the classification performance decreases as the distance threshold increases, as expected, since the further the vessels, the lower the quality of the spectral information. This is true except for the 5000m5000\,m threshold, which suddenly increases. This effect is due to the extreme data unbalance towards class 0 in that case, which leads to a close to zero $F_{1}^{C1}{}$ as can be seen in Fig. 10.c.
Regarding the impact of the spatial redundancy exploitation, we can observe a consistent tendency to get better results by using a higher number of sensors, which can be more clearly seen in Fig. 11, for which the data in Fig. 10 has been restricted to 500m500\,m, 1000m1000\,m and 1500m1500\,m thresholds. In the same way, the vertical axis range has been modified to facilitate the visualization.
Considering the results shown so far, it would seem that the classification task for the 500m500\,m distance threshold gets the highest performance, but it should be noted that it suffers from severe data unbalance (see Fig. 5), and its superiority is only true if we evaluate the $F_{1}^{G}{}$ or $F_{1}^{C1}$ scores. In fact, the best result selection should also consider the behavior of the class wise $F_{1}{}$-scores. In particular, as also shown in Fig. 11, the 500m500\,m threshold experiment $F_{1}^{C0}{}$s presents the lowest values. On the contrary, the behavior of the 1000m1000\,m threshold task exhibits a better balance between the $F_{1}^{C0}{}$ and $F_{1}^{C1}{}$ scores. So, from now on, we will focus on the 1000m1000\,m threshold task spatially integrating the feature vectors from $250$ sensors.
Figure 12: $F_{1}{}$-scores for 1000m1000\,m distance threshold as a function of: (a) temporal context duration ($S_{C}=\{10,30,50\}$ seconds long, T1000_V1_S[10,30,50]_ACH), (b) 10s10\,s temporal context with majority window length ($V_{C}=\{1,3,5\}$ windows, T1000_V[1,3,5]_S10_ACH), (c) 50s50\,s temporal context with majority window length ($V_{C}=\{1,3,5\}$ windows, T1000_V[1,3,5]_S50_ACH).
Figure 13: Global $F_{1}{}$-score as a function of the averaging type.
Fig. 12.a shows the $F_{1}{}$-scores for the 1000m1000\,m threshold with $10$, $30$ and $50$ second-long windows, with improvements for longer windows, specially in the $F_{1}^{C0}$ metric. Similarly, Figures 12.b and Fig. 12.c show the $F_{1}{}$-scores for the 1000m1000\,m threshold when using different majority voting window lengths ($1$, 33 or $5$) for temporal contexts of $10$ and $50$ seconds, respectively. From the results, when increasing the number of windows for majority voting, there are improvements for the smallest time span (10s10\,s), but the differences are small when the considered temporal context is increased (50s50\,s). This can be explained by the fact that the increased temporal context considers the same long temporal range than the increased voting scheme window length.
Figure 14: Classification task: Performance comparison between the best XGBoost and NN systems.
Finally, Fig. 13 shows the effect of spatial (ACH) vs. spatial+temporal averaging (ATI) for the 1000m1000\,m distance threshold, 50s50\,s temporal context, and $5$ windows of majority voting. The results show that the inclusion of temporal averaging is not useful when applied to strategies that already make use of the redundancy in temporal information (long temporal contexts and majority voting).
Therefore, considering the contributions of the algorithmic variations and aiming for a balanced performance across the $F_{1}{}$-scores, the final system we select for the production vessel-detection task is the one using a distance threshold of 1000m1000\,m (consistent with known criteria adopted by cable companies, as discussed in Section V-C), data from $250$ sensors, a $50$-second-length temporal context, majority voting with $5$ windows, and spatial averaging only.
We also ran the experiments using the simple NN-based system described in Section V-C. The observations regarding the effect of the algorithmic modifications are very similar to those previously discussed. However, the results are significantly worse than those obtained with the XGBoost system, as shown in Fig. 14. Future work will be devoted to a more exhaustive evaluation of variants and parameter tuning of both approaches.
VI-F Vessel Localization Experiments
As in the previous section, we will first describe the experiments using the XGBoost system, and then the comparison with the NN-based approach.
For the vessel localization task, we adopted the same experimental structure applied for vessel detection. The results for the baseline experiments are shown in Fig. 15.a. As expected, the distance-estimation MAE increases as the distance threshold increases. To better compare the impact of distance thresholds, Fig. 15.b provides the relative distance MAE, showing that the MAE ranges from approximately $12\%$ to $25\%$ of the distance threshold, decreasing as the threshold increases. This indicates that the system exhibits greater ambiguity when estimating closer vessel positions. A possible explanation is that closer vessels increase the likelihood of multiple vessels falling within the fiber’s detection range, thus interfering with the regression process.
As in the classification task, there is a consistent tendency to get better results when using a higher number of sensors, as shown in Fig. 15.c (again only using 500m500\,m, 1000m1000\,m and 1500m1500\,m thresholds).
Figure 15: Baseline MAE and rMAE scores (T[500,1000,1500,2000,3000,5000]_V1_S10_ACH): (a) Distance estimation MAE as a function of distance thresholds, (b) Relative MAE as a function of distance thresholds, (c) Distance estimation MAE as a function of the number of sensors used.
From now on, we will focus on the 1000m1000\,m distance threshold condition using $250$ sensors, as it was the one selected in the classification task.
Fig. 16.a shows the effect of considering longer temporal contexts. Again, exploiting longer temporal contexts improves the distance estimation results significantly.
Figure 16: Distance estimation MAE for the 1000m1000\,m distance threshold as a function of (a) temporal context duration ($S_{C}=\{10,30,50\}$ seconds long, T1000_V1_S[10,30,50]_ACH), (b) 10s10\,s temporal context with majority window length ($V_{C}=\{1,3,5\}$ windows, T1000_V[1,3,5]_S10_ACH), (c) 50s50\,s temporal context with majority window length ($V_{C}=\{1,3,5\}$ windows, T1000_V[1,3,5]_S50_ACH).
To assess the effect of majority voting strategies, Figures 16.b and 16.c show the MAE for the 1000m1000\,m threshold with 1, 3 or 5 majority voting windows, and using $10s$ or $50s$ time spans, respectively. Contrary to the findings in the classification task, we do observe significant improvements by increasing both the temporal context and the number of windows of the majority voting strategy in this regression task.
Finally, Fig. 17 shows the effect of spatial (ACH) vs spatial+temporal averaging (ATI). Contrary to the findings in the classification task, the inclusion of temporal averaging significantly deteriorates the estimated distances, which can be explained by considering that distance estimation is more severely affected by variations along the temporal span.
Figure 17: Distance estimation MAE for the 1000m1000\,m distance threshold, 50 seconds time span, 5 majority voting windows, evaluating spatial averaging (ACH) vs. spatial+temporal averaging (ACI) (T1000_V5_S50_A[CH,TI].
Therefore, considering the contributions of the algorithmic variations, the final system we would select for the production vessel distance-estimation task is the one using a distance threshold of 1000m1000\,m, data from 250 sensors, a 50 second long temporal context, majority voting with 5 windows, and spatial averaging only.
We also ran the experiments using the simple NN-based system, with similar observations regarding the effect of the algorithmic modifications. The achieved results are again significantly worse than those obtained with the XGBoost system, as shown in Fig. 18.
A supplementary material Web page showing videos on the system operating in detection and distance estimation modes is available at https://geintra-uah.org/psi/index.html and is also linked from the project GitHub repository.
Figure 18: Regression task: Performance comparison between the best XGBoost and NN systems.
VI-G Comparison with beamforming methods
To provide a reference baseline based on classical acoustic localization techniques, we implemented a beamforming-based distance estimation method following the approach presented by Paap et al. [paap2025leveragin], who conducted their study in the same geographical area as ours. In particular, we adopted the Steered Response Power (SRP) algorithm [dibiase2001robust], which is virtually identical to the implementation of the migration-based method used by Paap (sharing the same propagation assumptions and processing principles), but calculating power in each grid position instead of stacked absolute values of the DAS waveforms. The SRP algorithm was applied to the DAS strain data using the same 250 channels and 10-second temporal windows than in our experimental setup. To optimize the propagation model parameters, we applied Paap’s methodology [paap2025leveragin], scanning for the optimal sound speed by maximizing the coherency of back-projected waveforms. We tested speeds from 100m/s100\,m/s to 1900m/s1900\,m/s and determined a maximum response at 1750m/s1750\,m/s. This value is highly consistent with the 1700m/s1700\,m/s found by Paap et al. in the same region [paap2025leveragin] (also with a buried cable), and can be interpreted as evidence of significant acoustic energy transmission through the seabed top sediment layer, rather than through the water column.
For each time window, the SRP function was evaluated over a two-dimensional spatial grid with 20m20\,m resolution covering the vicinity of the monitored cable (a bounding box defined as 1000m1000\,m from the extreme cable positions of the used channels). The algorithm determines the most probable vessel position as the grid point maximizing the steered response power, from which we derived the corresponding distance to the cable. All results presented in this section correspond to vessels located within 1km1\,km of the cable.
Figure 19: Comparison of SRP beamforming and ML regression methods. (a) SRP map for a case where both methods perform well, particularly as the vessel is close to the cable. (b) SRP map for a vessel that is more difficult to detect; while both methods struggle, the ML regressor’s error is lower. (c) A time-series plot of a vessel crossing above the fiber, showing the distance estimations by SRP (red trace) and the ML regressor (gray trace) against the ground truth (green trace), illustrating a case where both methods track the vessel accurately. (d) A distance estimation time-series plot for a quieter or more challenging vessel, demonstrating a case where the ML regressor performs significantly better than the SRP method.
TABLE II: Distance estimation performance of SRP and ML-based regression for vessels within 1 km of the cable.
Method MAE [m] 95 % CI [m]
SRP Beamforming 225 [222.17, 227.97]
ML Regression (XGBoost) 171 [169.22, 172.83]
Table II summarizes the MAE results with the 95 % confidence intervals (CI) for both methods. Fig. 19.a shows a sample SRP map where the steered-response peak accurately identifies the vessel position, a scenario where both the beamformer and the ML regressor perform very well. In contrast, 19.b shows a more challenging case where the vessel is harder for both methods to detect, though the ML regressor ultimately yields a lower estimation error. The time-series plots in Figs. 19.c and 19.d compare the distance estimations over a time window centered on a vessel crossing, showing the true distance (estimated from AIS), the ML regressor prediction, and the SRP prediction. Fig. 19.c shows an example where both methods track the vessel effectively, whereas Fig. 19.d shows an example with a more challenging vessel where the ML regressor performs much better than the beamforming approach due to the presence of local maxima in the SRP map that do not correspond to the real vessel position.
Beamforming methods provide physically interpretable spatial maps, require no training data, and adapt easily to new deployments. However, their performance is highly sensitive to propagation model parameters that are difficult to control, such as the effective sound speed and the fiber-seabed coupling. Additionally, they require dense spatial grids that makes them computationally expensive, especially in long-cable deployments.
Figure 20: Distribution of absolute distance errors by (a) vessel size range, (b) vessel type (indicating also the average length (avg.len.) for the group type).
In contrast, the proposed ML-based approach avoids strong physical assumptions on the propagation model but depend on representative training data. Once trained, models validated on diverse cable segments can generalize well to new environments, providing faster and more accurate distance estimates by directly mapping signal features to vessel range, effectively exploiting spatiotemporal redundancies across sensors, and offering a more scalable and practical framework for continuous vessel monitoring.
VI-H Dependency on Vessel Size and Type
To assess the influence of vessel characteristics on localization accuracy, we analyzed the absolute distance error as a function of both vessel size and type. We evaluated the task with the 1000m1000\,m distance threshold, 250 channels and 10-second temporal windows (that achieved a MAE of 171m171\,m), with the results distribution boxplots shown in Fig. 20.
In general, larger vessels show the lowest errors, consistent with their expected stronger and more stable acoustic emissions. However, there are cases not fitting in this tendency, such as the Passenger Ships (mainly cruises), achieving higher error than the smaller Pilot Vessels, probably due to their expected lower radiated noise (being focused in passenger comfort).
As a conclusion, vessel size and type affect DAS-based localization performance, and its detailed study is open for future work, initially suggesting that class dependent models could further enhance system robustness and enable addressing vessel type identification tasks.
VII Conclusions and Future Work
Methods aimed at maritime surveillance in submarine cable deployments are crucial to prevent accidental damage or sabotage. This work represents, to the best of our knowledge, the first systematic study in open waters and realistic conditions to address vessel detection and localization specifically for submarine cable protection, at a much higher scale than previous works in the literature, establishing a foundational framework for leveraging existing underwater fiber infrastructure in security and monitoring applications.
Leveraging a 28km28\,km ocean‐bottom fiber in the Southern Bight of the North Sea and ten days of continuous DAS and AIS data monitoring, we developed and evaluated a systematic, data-driven processing pipeline that includes:
  • A data-driven spectral feature extraction based on logarithmically-spaced frequency‐band energy measurements to capture vessel acoustic signatures.
  • Two machine learning frameworks for (i) vessel detection posed as a binary classification problem (vessel closer of further than a predefined threshold distance) and (ii) vessel distance estimation posed as a regression problem.
  • A rigorous evaluation over more than $74000$ data frames, adopting a $k$-fold cross-validation strategy that systematically avoids data leakage and increases data variability.
  • A systematic exploitation of spatial (up to 250 channels over 2.5km2.5\,km) and temporal redundancies to enhance robustness.
For the selected 1000m1000\,m threshold distance, our best performing vessel detection model achieved an overall $F_{1}{}$-score of over $90\%$, with balanced class-wise performance; and for vessel localization, the selected regression system achieved a mean absolute error of 141m141\,m, demonstrating reliable detection and distance estimation under diverse environmental and traffic conditions. We also carried out a comparison with a beamforming-based distance estimation method, showing that the proposed ML approach achieves lower localization errors and offers a more robust and computationally efficient alternative. These results highlight DAS as a practical alternative to conventional maritime surveillance methods, allowing for continuous, all-weather monitoring without reliance on cooperative systems such as AIS.
The release of the full dataset used in this work will enable other researchers to easily reproduce our machine learning experiments, compare new algorithms, and extend vessel detection efforts in other submarine cable-related contexts. Our field validation focused on vessel detection and localization for submarine cable protection represents the most extensive DAS-based vessel monitoring study to date, opening the path for several research areas to be explored as future work. Our first task will be devoted to applying the methodology to additional submarine cable deployments and emerging public DAS datasets [ais-data-ieee-dataport-1308-8605-25], to assess generalization across different bathymetries, seabed couplings, vessel types (proved to have an impact in localization accuracy), traffic profiles, and weather conditions. Beyond binary detection and range estimation, we will also address ML tasks to infer vessel speed, size, and type, as well as multi-target tracking and precise localization, by integrating richer AIS metadata, advanced signal processing and Deep Learning strategies. Exploiting meteorological data and bathymetric variables will be considered, either into feature sets or as model conditioning factors, given their significant influence on DAS submarine data acquisition. Finally, for real-world deployments, we will need to adopt efficient, low-latency monitoring pipelines such as the recent proposal in [biondi2025real].